1) Convert 1 millisecond to hours Before starting, complete the unit map __millisecond___  _________  _________  __hours___


2) Convert 2500 mg to kg (use one unit map or the other, don't need both

3) Convert 8.7 cm to mm (use one unit map or the other, don't need both) _________  _________  _________ or _________  _________

4) Convert 8 yards to meters (1 ft = 12 in, 2.54 cm = 1 in) __yards___  feet  _________  _________  meters _   
Answer #1

Q.) I minisecond to hours . The I millisecond = 10-3 second = 10-3 minutes = 1073 Preurs => 2.78x10-7 hours 60x60 Q.2 2500mg